Evaluate P(-1/2) if P(x) = 2x4 + x3 + 12
I don't really know what this equation is wanting me to do....????

Answer by uknouluvme89(30)

Evaluate P(-1/2) if P(x) = 2x4 + x3 + 12
this equation is giving you the value -1/2 for x and is asking you to subsitute it into the equation.
P(-1/2) = 2x4 + x3 + 12
P(-1/2) = 2(.0625)-.125 + 12
P(-1/2) = .125 -.125 + 12
P(-1/2) = 12